OpenStudy (anonymous):

simplify using the laws of exponents (a^4b)^5

hartnn (hartnn):

\(\huge (xy)^n =x^n.y^n\) so take x=a^4 and y=b, n=5 and tell me what u get ?

OpenStudy (anonymous):

(4^b)^54=a^4*b^4

OpenStudy (anonymous):

that is not supposed to be 54

hartnn (hartnn):

no... \(\huge (a^4b)^5=(a^4)^5b^5\) got that ^ ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

now , \(\huge (x^m)^n=x^{mn}\) so whats \((a^4)^5=....?\)

OpenStudy (anonymous):

16^5

hartnn (hartnn):

what ? where does 16 come from ??

OpenStudy (anonymous):

20^5 sorry

hartnn (hartnn):

where did 'a' go ?

OpenStudy (anonymous):

20a^5

hartnn (hartnn):

nopes, the exponents get multiplied. \((a^4)^5=a^{5*4}=a^{20}\) got that ^ ?

OpenStudy (anonymous):

yes got that

hartnn (hartnn):

so whats the final answer ?

OpenStudy (anonymous):

a^20

OpenStudy (anonymous):

sorry im getting confused would it be a fraction form is so it would be 1/20

hartnn (hartnn):

no..its just a^20

OpenStudy (anonymous):

ok ty

hartnn (hartnn):

\(\huge (a^4b)^5=(a^4)^5b^5=a^{20}b^5\) is your final answer.